# Variance-Reducing Couplings for Random Features: Perspectives from Optimal Transport

### v1

@@ Abstract
Random features (RFs) are a popular technique to scale up kernel methods in machine learning, replacing exact kernel evaluations with stochastic Monte Carlo estimates. They underpin models as diverse as efficient transformers (by approximating attention) to sparse spectrum Gaussian processes (by approximating the covariance function). Efficiency can be further improved by speeding up the convergence of these estimates: a variance reduction problem. We tackle this through the unifying framework of optimal transport, using theoretical insights and numerical algorithms to develop novel, high-performing RF couplings for kernels defined on Euclidean and discrete input spaces. They enjoy concrete theoretical performance guarantees and sometimes provide strong empirical downstream gains, including for scalable approximate inference on graphs. We reach surprising conclusions about the benefits and limitations of variance reduction as a paradigm.
@@ 1 Introduction
Random features for kernel approximation. Though very effective on small datasets, kernel methods suffer from poor scalability. The need to materialise and invert the Gram matrix $\mathbf{K}\coloneqq[k(\bm{x}_{i},\bm{x}_{j})]_{i=1}^{N}$ leads to a time complexity cubic in the size of the dataset $N$. Substantial research has been dedicated to improving scalability by approximating this matrix, a prominent example being random features (RFs) (Rahimi and Recht, 2007, 2008; Avron et al., 2017b; Liu et al., 2022). These randomised mappings $\phi:\mathbb{R}^{d}\to\mathbb{R}^{s}$ construct low-dimensional or sparse feature vectors that satisfy
@@
The features $\phi(\bm{x})$ are typically constructed using ensembles of random frequencies $\{\bm{\omega}\}_{i=1}^{m}$ drawn from some known distribution $\mathcal{\eta}$. The choice of $\eta$ and manner in which they are combined to construct $\phi(\bm{x})$ depends on the particular input space $\mathcal{X}$ and kernel function $k$ being approximated. The set of RFs $\{\phi(\bm{x}_{i}\}_{i=1}^{N}$ can then be used to construct a low-rank or sparse approximation of the Gram matrix, providing substantial space and time complexity savings. RFs exist for a variety of kernels, including for continuous and discrete input spaces (Dasgupta et al., 2010; Johnson, 1984; Choromanski et al., 2020; Goemans and Williamson, 2001; Rahimi and Recht, 2007; Choromanski, 2023).
@@
Variance reduction for RFs. Eq. 1 can be understood as a Monte Carlo (MC) estimate of $k$. In practical applications, it is often found that this estimate converges slowly. This can be addressed by taking many features $m$, but this undermines the efficiency gains of RFs. Therefore, substantial effort has been dedicated to reducing the variance of the kernel estimates. Variance reduction methods include quasi-Monte Carlo (QMC) (Dick et al., 2013; Yang et al., 2014a), common random numbers (CRNs) (Glasserman and Yao, 1992), antithetic variates (Hammersley and Morton, 1956) and structured Monte Carlo (SMC) (Yu et al., 2016), all of which are well-studied in computational statistics. These techniques all ultimately work by replacing i.i.d. frequencies $\{\bm{\omega}_{i}\}_{i=1}^{m}$ by a dependent ensemble, with the dependencies between samples designed to improve RF convergence.
@@
Limitations of previous techniques. The best choice of dependencies between $\{\bm{\omega}_{i}\}_{i=1}^{m}$ is an active research area. Though straightforward to apply, standard QMC techniques are suboptimal. They are based on fixed ‘low-discrepancy sequences’ so typically do not incorporate information about the particular kernel function $k$ being approximated. Empirical performance may be poor and theoretical guarantees lacking in the low-sample, high-dimensional regime (Rowland et al., 2018; Morokoff and Caflisch, 1995) – which is precisely where RFs are most important. On the other hand, hand-crafted SMC dependencies – which impose strict geometrical conditions like orthogonality between frequencies – tend to fare better (Yu et al., 2016). But they are still difficult to design, theoretical guarantees are hard-won and optimality is not guaranteed. RFs for estimating kernels defined on discrete spaces like the nodes of a graph have only recently been developed (Choromanski, 2023; Tripp et al., 2024), so here very few effective variance reduction techniques have even been proposed.
@@
Optimal transport. To address these shortcomings, we propose to frame variance reduction as optimal transport: an active research area of applied mathematics that studies how to move (probability) mass between distributions as efficiently as possible (Villani et al., 2009). This fruitful perspective equips us with proof techniques and numerical tools to identify the best possible dependencies between samples, giving lower kernel estimator variance compared to previous approaches. We use these insights to develop better couplings for RFs in Euclidean and discrete spaces.
Our contributions. We summarise our key contributions below.
@@
    For random Fourier features (RFFs) and random Laplace features (RLFs), we exactly solve the OT problem for the norms of $m=2$ orthogonal frequencies, finding a novel coupling with the lowest possible variance (Sec. 3). We use a copula as a numerical OT solver to consider arbitrary $m$.
@@
    For graph random features (GRFs), we couple the lengths of random walks by solving an OT bipartite matching, outcompeting previous algorithms for estimating graph node kernels (Sec. 4).
@@
    We test our algorithms on UCI datasets and real-world graphs, verifying that OT couplings substantially reduce kernel estimator variance (Secs 3 and 4). We show that this sometimes translates to much better performance in downstream tasks (in particular, approximate inference with scalable graph-based Gaussian processes), but also reach surprising conclusions about the limitations of variance reduction for RFs in machine learning.
@@
All proofs are saved for the Appendices, but are also sketched in the main body where space allows. We will make source code available if accepted.
2 Background
@@
From kernel estimation to optimal transport (OT). Considering the expression in Eq. 1, define the kernel estimator $\widehat{k}(\bm{x}_{i},\bm{x}_{j})\coloneqq\phi(\bm{x}_{i})^{\top}\phi(\bm{x}_{j})$. Recall that $\phi(\cdot)$ is computed using random frequencies $\{\bm{\omega}_{i}\}_{i=1}^{m}$, with the space in which they live, distribution from which they are drawn, and manner in which they are combined dependent on the particular kernel being approximated. The estimator is unbiased provided each frequency $\bm{\omega}_{i}$ obeys some known, fixed marginal distribution $\eta$. Importantly, independence of $\{\bm{\omega}_{i}\}_{i=1}^{m}$ is not a requirement: any joint distribution with marginals $\eta$ gives an unbiased estimator. We refer to the set of such joint distributions as couplings.
The choice of coupling between the $m$ frequencies determines the variance of the estimator. It is straightforward to see that finding the coupling $\mu^{*}$ that minimises this variance is equivalent to solving
$\mu^{*}=\textrm{arg inf}_{\mu\in\Lambda_{m}(\eta)}\left[\mathbb{E}_{\bm{\omega}_{1:m}\sim\mu}c(\bm{\omega}_{1:m})\right]$ (2)
where we defined the cost function
$c(\bm{\omega}_{1:m})\coloneqq\left(\phi(\bm{x})^{\top}\phi(\bm{y})\right)^{2}.$ (3)
@@
$\Lambda_{m}(\eta)$ denotes the set of couplings over $m$ random variables with marginal measures $\eta$. This is precisely the Kantorovich formulation of a multi-marginal optimal transport problem, with cost function $c$ (see e.g. Eq. 4 of the seminal OT text of Villani (2021)). The relationship between variance reduction and optimal transport was also noted by Rowland et al. (2018), but in the different setting of estimating the expected value of functions drawn from Gaussian processes.
(Approximately) solving the OT problem. For many popular RF constructions, the OT problem in Eq. 2 is intractable. The cost function and corresponding optimal coupling will in general depend on the choice of RF map and data distribution. A further practical consideration is that, to be useful, $\mu^{*}$ must be easy to sample from; even if it guarantees lower variance, an expensive sampling mechanism will obviate any efficiency gains from a coupling. However, we will see that for certain RFs under particular assumptions we can solve the OT problem exactly, obtaining computationally lightweight couplings that provide the smallest possible kernel estimator variance. Even if the OT problem is intractable, we can often leverage numerical methods to approximately solve Eq. 2. The formulation of these methods will depend on the input domain of the RF being considered.
Remainder of the manuscript. We will demonstrate how insights and algorithms from OT can be used to improve the convergence of RFFs and RLFs (Sec. 3) and GRFs (Sec. 4). These have input domains $\mathcal{X}=\mathbb{R}^{d}$ ($d$-dimensional Euclidean space) and $\mathcal{X}=\mathcal{N}$ (the set of graph nodes), respectively.
@@ 3 Random Fourier features and random Laplace features
where $\odot$ denotes concatenation. These provide an unbiased estimate if the frequencies $\{\bm{\omega}_{i}\}_{i=1}^{m}$ are marginally Gaussian, $\bm{\omega}_{i}\sim\mathcal{N}(0,\mathbf{I}_{d})$. They are widely used for scaling kernel methods such as Gaussian processes (GPs; Williams and Rasmussen, 2006) and support vector machines (SVMs; Scholkopf and Smola, 2018). The time complexity of computing the exact posterior of a GP is $\mathcal{O}(N^{3}),$ where $N$ is the number of datapoints. Using RFFs, one can approximate the posterior with $m\ll N$ features, reducing this cost to $\mathcal{O}(Nm^{2})$. Beyond RFFs, $k(\bm{x}_{i},\bm{x}_{j})$ can also be approximated using random Laplace features (RLFs) (Yang et al., 2014b),
@@
where again $\bm{\omega}_{i}\sim\mathcal{N}(0,\mathbf{I}_{d}).$ Unlike RFFs, RLFs give strictly positive kernel estimates. This makes them better suited to estimating attention in linear-attention transformers (Choromanski et al., 2020), where negative estimates can cause training instabilities. Using $m$ RLFs to write a low-rank decomposition of the attention matrix with $N$ $d$-dimensional tokens, one can reduce the computational cost of transformers from $\mathcal{O}(N^{2}+Nd)$ down to $\mathcal{O}(Nmd)$ with low performance loss.
Orthogonal random features. A common variance reduction technique for both RFFs and RLFs is the orthogonality trick (Yu et al., 2016; Rowland et al., 2018; Reid et al., 2023; Choromanski et al., 2018). Exploiting the isotropy of $\mathcal{N}(0,\mathbf{I}_{d})$, it is possible to constrain the frequency vectors $\{\bm{\omega}_{i}\}_{i=1}^{m}$ to be exactly orthogonal whilst preserving their marginal distributions. This is found to reduce the kernel estimator variance and improve performance in downstream tasks. Whilst this technique couples the directions of the random frequencies $\{\widehat{\bm{\omega}}_{i}\}_{i=1}^{m}$, their norms $\{\omega_{i}\}_{i=1}^{m}$ (with $\omega_{i}\coloneqq\|\bm{\omega}_{i}\|$) are left independent so the coupling is suboptimal. We will now use tools from OT to find a coupling between frequency vector norms that provides strictly lower kernel estimator variance.
@@ 3.1 Solving the OT problem for maximal variance reduction
Consider an ensemble of $m$ orthogonal random frequency directions $\{\widehat{\bm{\omega}}_{i}\}_{i=1}^{m}$, jointly randomly rotated so they are marginally isotropic. Recall that our task is to couple their norms $\{\omega_{i}\}_{i=1}^{m}$ to suppress the RFF and RLF kernel estimator variance. The marginal distribution of each $\omega_{i}$ must be $\chi_{d}$ (a Chi distribution with $d$ degrees of freedom) to ensure that each $\bm{\omega}_{i}$ is marginally Gaussian. It is simple to extend recently-derived results (Reid et al., 2023) to compute the OT cost functions.
@@
$\mu^{*}=\arg\inf_{\mu\in\Lambda_{m}(\chi_{d})}\left[\mathbb{E}_{\omega_{1:m}\sim\mu}c(\omega_{1:m})\right],\quad\textrm{where}$ (6)
@@
with $z\coloneqq\|\bm{x}-\bm{y}\|_{2}$ and $v\coloneqq\|\bm{x}+\bm{y}\|_{2}$.
This is a challenging multi-marginal OT problem. However, remarkably, we can solve it exactly, under mild asymptotic assumptions for RFFs, when $m=2$. The following result is novel.
@@
Denote by $F_{\chi_{d}}(\cdot)$ the cumulative distribution function (CDF) of $\chi_{d}$. Considering $m=2$ orthogonal frequencies with norms $(\omega_{1},\omega_{2})$, the OT problem in Eq. 6 is solved by the negative monotone coupling
$F_{\chi_{d}}(\omega_{1})+F_{\chi_{d}}(\omega_{2})=1,$ (8)
when $\frac{z^{2}}{d}\to 0$ for RFFs and in full generality for RLFs.
Proof sketch. We defer a full proof of this important result to App. A.2, but provide a sketch for the interested reader. It is well known that optimal transport plans satisfy a property called ‘$c$-monotonicity’, which specifies how the support of the optimal coupling depends on the cost function. For $c_{\textrm{RFF}}$ and $c_{\textrm{RLF}}$, we can show that this is commensurate with negative monotonicity (Eq. 8), though only asymptotically for the former. ∎
Given $m=d$ orthogonal frequencies, it is possible to partition the ensemble into $\lfloor\frac{d}{2}\rfloor$ orthogonal pairs, with one remaining frequency if $d$ is odd. For every such pair, we can impose negative monotone coupling (Eq. 8). We refer to such dependent ensembles as pairwise norm-coupled (PNC).
@@
We refer to RFs as pairwise norm-coupled (PNC) if $d$ orthogonal frequencies $\{\bm{\omega}_{i}\}_{i=1}^{d}$ are arranged in $\lfloor\frac{d}{2}\rfloor$ pairs, each of which exhibits negative montone coupling so that $F_{\chi_{d}}(\omega_{1})+F_{\chi_{d}}(\omega_{2})=1$. Different pairs are independent.
To reduce the variance further, one can of course take multiple independent PNC ensembles. A corollary of Thm. 3.2 is as follows.
Corollary 3.4 (Supriority of pairwise norm-coupled RFs).
The kernel estimator variance with pairwise norm-coupled RFs is strictly lower than orthogonal RFs with independent norms, provided $\frac{z^{2}}{d}\to 0$ for RFFs and in full generality for RLFs.
Negative monotone coupling differs from OT plans usually seen in machine learning; it is a space-filling coupling that seeks long transport plans that give diverse samples. However, it is a popular heuristic technique for variance reduction via common random numbers (CRNs) in computational statistics (Glasserman and Yao, 1992). This is (to our knowledge) the first result applying it to improving the convergence of orthogonal RFs, and the first corresponding guarantees for variance reduction. We make one further theoretical contribution regarding coupled RLFs.
@@
For RLFs with $m=2$ frequencies whose respective orientations $(\widehat{\bm{\omega}}_{1},\widehat{\bm{\omega}}_{2})$ are unconstrained, the variance reduction OT problem is solved by conditioning that $\bm{\omega}_{1}=-\bm{\omega}_{2}$ almost surely (that is, opposite directions and equal norms).
This coupling is popularly known as antithetic sampling (Hammersley and Morton, 1956). Thm. 3.5 shows that, given a PNC ensemble $\{\bm{\omega}_{i}\}_{i=1}^{d}$, we can obtain further variance reduction by augmenting it to {±$\{\pm\bm{\omega}_{i}\}_{i=1}^{d}$. Antithetic sampling is also a common (though often heuristically motivated) variance reduction strategy used e.g. when estimating attention in Performers (Choromanski et al., 2020). We can reinterpret its effectiveness through the lens of OT. Note that Thm. 3.5 does not hold for RFFs.

### v2

@@ Abstract
Random features (RFs) are a popular technique to scale up kernel methods in machine learning, replacing exact kernel evaluations with stochastic Monte Carlo estimates. They underpin models as diverse as efficient transformers (by approximating attention) to sparse spectrum Gaussian processes (by approximating the covariance function). Efficiency can be further improved by speeding up the convergence of these estimates: a variance reduction problem. We tackle this through the unifying lens of optimal transport, finding couplings to improve RFs defined on both Euclidean and discrete input spaces. They enjoy theoretical guarantees and sometimes provide strong downstream gains, including for scalable approximate inference on graphs. We reach surprising conclusions about the benefits and limitations of variance reduction as a paradigm, showing that other properties of the coupling should be optimised for attention estimation in efficient transformers.
@@ 1 Introduction
Random features for kernel approximation. Though very effective on small datasets, kernel methods suffer from poor scalability. The need to materialise and invert the Gram matrix $\mathbf{K}\coloneqq[k(\bm{x}_{i},\bm{x}_{j})]_{i=1}^{N}$ leads to a time complexity cubic in the size of the dataset $N$. Substantial research has been dedicated to improving scalability by approximating this matrix, a prominent example being random features (RFs) (Rahimi and Recht, 2007; 2008; Avron et al., 2017b; Liu et al., 2022). These randomised mappings $\phi:\mathbb{R}^{d}\to\mathbb{R}^{s}$ construct low-dimensional or sparse feature vectors that satisfy
@@
The expectation $\mathbb{E}$ is taken over an ensemble of random frequencies $\{\bm{\omega}\}_{i=1}^{m}$ drawn from a distribution $\mathcal{\eta}$. The space in which $\{\bm{\omega}\}_{i=1}^{m}$ live and manner in which they are combined to construct $\phi(\bm{x})$ depends on the particular input space $\mathcal{X}$ and kernel function $k$ being approximated. This paper will consider several examples. The set of RFs $\{\phi(\bm{x}_{i})\}_{i=1}^{N}$ can be used to construct a low-rank or sparse approximation of the Gram matrix, providing substantial space and time complexity savings. RFs exist for a variety of kernels, including for continuous and discrete input spaces (Dasgupta et al., 2010; Johnson, 1984; Choromanski et al., 2020; Goemans and Williamson, 2001; Rahimi and Recht, 2007; Choromanski, 2023; Tripp et al., 2024).
@@
Variance reduction for RFs. Eq. 1 can be understood as a Monte Carlo (MC) estimate of $k$. In applications, it is often found that this estimate converges slowly. This can be addressed by taking many samples $m$, but this undermines the efficiency gains of RFs. Therefore, substantial effort has been dedicated to reducing the variance of the kernel estimates. Variance reduction methods include quasi-Monte Carlo (QMC; Dick et al., 2013; Yang et al., 2014a), common random numbers (CRNs; Glasserman and Yao, 1992), antithetic variates (Hammersley and Morton, 1956) and structured Monte Carlo (SMC; Yu et al., 2016). These techniques work by replacing i.i.d. frequencies $\{\bm{\omega}_{i}\}_{i=1}^{m}$ by a dependent ensemble, with the sample dependencies designed to improve RF convergence.
@@
Limitations of previous techniques. The best choice of dependencies between $\{\bm{\omega}_{i}\}_{i=1}^{m}$ is an active research area. Though straightforward to apply, standard QMC techniques are suboptimal. They are based on hard-coded ‘low-discrepancy sequences’ so typically do not incorporate information about the particular kernel $k$ being approximated. Empirical performance may be poor and theoretical guarantees lacking in the low-sample, high-dimensional regime (Rowland et al., 2018; Morokoff and Caflisch, 1995), which is precisely where RFs are most important. On the other hand, hand-crafted SMC dependencies, which impose strict geometrical conditions like orthogonality between frequencies, tend to fare better (Yu et al., 2016). But they are difficult to design, theoretical guarantees are hard-won and optimality is not guaranteed. RFs for estimating kernels defined on discrete spaces like the nodes of a graph have only recently been developed (Choromanski, 2023; Tripp et al., 2024), so here very few effective variance reduction techniques have even been proposed. This paper asks: can we devise a principled framework for coupling RFs, providing variance reduction across basis functions and input domains, including with very few samples?
@@
Optimal transport. To answer this, we propose to frame variance reduction as optimal transport (OT): an active research area of applied mathematics that studies how to move (probability) mass between distributions as efficiently as possible (Villani et al., 2009). This novel perspective equips us with proof techniques and numerical tools to identify the best possible dependencies between samples, giving lower kernel estimator variance compared to previous approaches. OT allows us to improve couplings for RFs in both Euclidean and discrete spaces, including with different basis functions. To our knowledge, this has never before been achieved in the same paper.
Our contributions. This work presents unifying strategies to reduce the variance of random features.
@@
    For random Fourier features (RFFs) and random Laplace features (RLFs), we exactly solve the OT problem for the norms of $m=2$ orthogonal frequencies (Sec. 3). We introduce pairwise norm-coupling, which guarantees lower variance for arbitrary $m$.
@@
    For graph random features (GRFs), we couple the lengths of random walks by finding a bipartite matching between the quantiles of the marginal distributions (Sec. 4). This is the first time a coupling between random walks has been optimised using data, beating hard-coded algorithms.
@@
    We test our algorithms on UCI datasets and real-world graphs, verifying that OT couplings substantially reduce kernel estimator variance (Secs 3 and 4). We show that this sometimes translates to much better performance in downstream tasks, including for approximate inference with scalable graph-based Gaussian processes. However, we also reach surprising conclusions about the limitations of variance reduction for RFs, including for efficient transformers.
@@
All proofs are saved for the Appendices, but are also sketched in the main body where space allows.
2 Preliminaries
@@
From kernel estimation to optimal transport (OT). Define the kernel estimator $\widehat{k}(\bm{x}_{i},\bm{x}_{j})\coloneqq\phi(\bm{x}_{i})^{\top}\phi(\bm{x}_{j})$. Recall that $\phi(\cdot)$ is computed using random frequencies $\{\bm{\omega}_{i}\}_{i=1}^{m}$, with the space in which they live, distribution from which they are drawn, and manner in which they are combined dependent on the particular kernel being approximated. The estimator is unbiased provided each frequency $\bm{\omega}_{i}$ obeys some marginal distribution $\eta$. Importantly, independence of $\{\bm{\omega}_{i}\}_{i=1}^{m}$ is not required: any joint distribution with marginals $\eta$ gives an unbiased estimator. We refer to the set of such joint distributions as couplings.
The coupling between the frequencies determines the estimator variance. We want to solve:
$\textrm{minimise }\mathcal{I}(\mu)=\mathbb{E}_{\bm{\omega}_{1:m}\sim\mu}c(\bm{\omega}_{1:m})\quad\textrm{ for }\quad\mu\in\Lambda_{m}(\eta),$ (2)
@@
where we defined the cost function $c(\bm{\omega}_{1:m})\coloneqq\left(\phi(\bm{x})^{\top}\phi(\bm{y})\right)^{2}$ and $\Lambda_{m}(\eta)$ denotes the set of couplings of $m$ random variables with marginal measures $\eta$. This is precisely the Kantorovich formulation of a multi-marginal OT problem (see Eq. 4 of the seminal OT text of Villani (2021)). We will generally consider cost functions where the minimiser exists and we want to find efficient new MC couplings, so the task is to find the optimal coupling $\mu^{*}=\arg\min_{\mu\in\Lambda_{m}(\chi_{d})}\left[\mathbb{E}_{\omega_{1:m}\sim\mu}c(\omega_{1:m})\right]$ with the smallest estimator variance. The relationship between variance reduction and OT was also noted by Rowland et al. (2018) in a different context.
(Approximately) solving the OT problem. The formulation of Eq. 2 depends on the particular RF mechanism and kernel being approximated. We will show that one can solve it exactly for RFFs and RLFs (Sec. 3) and approximately for GRFs (Sec. 4), which have input domains $\mathcal{X}=\mathbb{R}^{d}$ ($d$-dimensional Euclidean space) and $\mathcal{X}=\mathcal{N}$ (the set of graph nodes) respectively. This gives new couplings with lower RF variance than previous algorithms.
@@ 3 Random Fourier features and random Laplace features
where $\odot$ denotes concatenation. These provide an unbiased estimate if the frequencies $\{\bm{\omega}_{i}\}_{i=1}^{m}$ are marginally Gaussian, $\bm{\omega}_{i}\sim\mathcal{N}(0,\mathbf{I}_{d})$. RFFs are widely used for scaling kernel methods such as Gaussian processes (GPs; Williams and Rasmussen, 2006) and support vector machines (SVMs; Scholkopf and Smola, 2018). The time complexity of computing the exact posterior of a GP is $\mathcal{O}(N^{3}),$ where $N$ is the number of datapoints. Using RFFs, one can approximate the posterior with $m\ll N$ features, reducing this cost to $\mathcal{O}(Nm^{2})$. Changing basis functions, $k(\bm{x}_{i},\bm{x}_{j})$ can also be approximated using random Laplace features (RLFs) (Yang et al., 2014b),
@@
where again $\bm{\omega}_{i}\sim\mathcal{N}(0,\mathbf{I}_{d}).$ Unlike RFFs, RLFs guarantee positive kernel estimates. This makes them better suited to approximating attention in efficient transformers (Choromanski et al., 2020), where negative estimates cause training instabilities. Using $m$ RLFs to get a low-rank decomposition of attention with $N$ $d$-dimensional tokens, one can reduce the time complexity of transformers from $\mathcal{O}(N^{2}+Nd)$ to $\mathcal{O}(Nmd)$ with low performance loss.
Orthogonal random features. A common variance reduction technique for both RFFs and RLFs is the orthogonality trick (Yu et al., 2016; Rowland et al., 2018; Reid et al., 2023; Choromanski et al., 2018). Exploiting the isotropy of $\mathcal{N}(0,\mathbf{I}_{d})$, one can constrain the frequency vectors $\{\bm{\omega}_{i}\}_{i=1}^{m}$ to be exactly orthogonal whilst preserving their marginal distributions. This is found to reduce the kernel estimator variance and improve performance in downstream tasks. Whilst this technique couples the directions of the random frequencies $\{\widehat{\bm{\omega}}_{i}\}_{i=1}^{m}$, their norms $\{\omega_{i}\}_{i=1}^{m}$ (with $\omega_{i}\coloneqq\|\bm{\omega}_{i}\|_{2}$) are left independent so the coupling is suboptimal. By solving an OT problem, we will show how coupling the norms can further reduce estimator variance.
@@ 3.1 Solving the OT problem for maximal variance reduction
Consider an ensemble of $m$ orthogonal random frequency directions $\{\widehat{\bm{\omega}}_{i}\}_{i=1}^{m}$, jointly randomly rotated so they are marginally isotropic. Our task is to couple their norms $\{\omega_{i}\}_{i=1}^{m}$ to suppress the RFF and RLF kernel estimator variance. The marginal distribution of each $\omega_{i}$ must be $\chi_{d}$ (a Chi distribution with $d$ degrees of freedom) to ensure that each $\bm{\omega}_{i}$ is marginally Gaussian. We can extend recent results by Reid et al. (2023) to compute the OT cost functions.
@@
$\mu^{*}=\arg\min_{\mu\in\Lambda_{m}(\chi_{d})}\left[\mathbb{E}_{\omega_{1:m}\sim\mu}c(\omega_{1:m})\right],\quad\textrm{where}$ (5)
@@
with $z\coloneqq\|\bm{x}-\bm{y}\|_{2}$ and $v\coloneqq\|\bm{x}+\bm{y}\|_{2}$. $\Gamma$ is the gamma function.
This is a tough multi-marginal OT problem. However, remarkably, we can solve it exactly, under mild asymptotic assumptions for RFFs, when $m=2$. The following result is novel.
@@
Denote by $F_{\chi_{d}}(\cdot)$ the cumulative distribution function (CDF) of $\chi_{d}$. Consider $m=2$ orthogonal frequencies with norms $(\omega_{1},\omega_{2})$. For RLFs, the OT problem in Eq. 5 is solved by the negative monotone coupling
$F_{\chi_{d}}(\omega_{1})+F_{\chi_{d}}(\omega_{2})=1.$ (7)
For RFFs, Eq. 7 ensures lower cost than any other coupling, provided $z$ is sufficiently small.
Proof sketch. We defer a full proof of this important result to App. A.2; here is a brief sketch. OT plans satisfy a property called ‘$c$-monotonicity’, which specifies how the support of the optimal coupling depends on the cost function. For RLFs, $c_{\textrm{RLF}}$ immediately implies negative monotonicity (Eq. 7). For RFFs, this is only true for the first nontrivial term in $z$. By bounding the contribution from the remaining terms, one can show that Eq. 7 still guarantees lower variance than any other coupling if $z$ is small enough. Specifically, letting $\mu_{\textrm{NM}}$ denote the negative monotone coupling, for any other coupling $\mu^{\prime}\in\Lambda_{2}(\eta)\backslash\left\{\mu_{\textrm{NM}}\right\}$ there exists some constant $\delta(\mu^{\prime})>0$ such that $\mathcal{I}(\mu_{\textrm{NM}})<\mathcal{I}(\mu^{\prime})$ for all $z<\delta$ (Lemma A.6). ∎
Given $m=d$ orthogonal frequencies, one can partition the ensemble into $\lfloor\frac{d}{2}\rfloor$ orthogonal pairs, with one remaining frequency if $d$ is odd. For every pair, one can impose negative monotone coupling (Eq. 7). We refer to such ensembles as pairwise norm-coupled (PNC).
@@
RFs are pairwise norm-coupled (PNC) if $d$ orthogonal frequencies $\{\bm{\omega}_{i}\}_{i=1}^{d}$ are arranged in $\lfloor\frac{d}{2}\rfloor$ pairs, each of which is negative montone-coupled so that $F_{\chi_{d}}(\omega_{1})+F_{\chi_{d}}(\omega_{2})=1$. Different pairs are independent.
PNC is no more expensive than i.i.d. norms. To reduce the variance further, one can take multiple independent PNC ensembles. An important corollary of Thm. 3.2 is as follows.
Corollary 3.4 (Superiority of pairwise norm-coupled RFs).
For any $m$, the variance of pairwise norm-coupled RFs is guaranteed to be lower than orthogonal RFs with independent norms, in full generality for RLFs and provided $z$ is small enough for RFFs.
Negative monotone coupling differs from OT plans usually seen in machine learning; it is a space-filling coupling that seeks long transport plans that give diverse samples. However, it is a popular heuristic technique for variance reduction via common random numbers (CRNs) in computational statistics (Glasserman and Yao, 1992). To our knowledge, this is the first result applying it to improving the convergence of orthogonal RFs, and the first corresponding guarantees for variance reduction. We make one further theoretical contribution for RLFs.
@@
For RLFs with $m=2$ frequencies whose respective orientations $(\widehat{\bm{\omega}}_{1},\widehat{\bm{\omega}}_{2})$ are unconstrained, variance is minimised by conditioning that $\bm{\omega}_{1}=-\bm{\omega}_{2}$ almost surely (that is, opposite directions and equal norms).
This coupling is known as antithetic sampling (Hammersley and Morton, 1956). Thm. 3.5 shows that, given a PNC ensemble $\{\bm{\omega}_{i}\}_{i=1}^{d}$, we can obtain further variance reduction by augmenting it to {±$\{\pm\bm{\omega}_{i}\}_{i=1}^{d}$. Antithetic sampling is also a common (though often heuristically motivated) variance reduction strategy used e.g. when estimating attention in Performers (Choromanski et al., 2020). We can reinterpret its effectiveness as an OT coupling.